FEL Upgrade Project Weekly Brief
January 14-18, 2002
This week we held our 13th Laser Processing Consortium/FEL Users Workshop on Jan. 17-18th. By all accounts it was a useful and informative meeting with 121 registrants. The meeting program is posted on our website. The usual informal proceedings will be made available to the registrants and other interested parties when we have collected all the presentations.

We used the LPC Workshop as an opportunity to have concurrent meetings of the FEL Program Advisory Committee and the Industrial Advisory Board. The advisory bodies provide valuable criticism and counsel on our FEL program and plans. We expect to receive their formal reports in the next few weeks.

On Friday, Jan. 18, we were asked by the Joint Technology Office to send a brief summary report on our FY01 funded activities on laser technology and we responded by the end of the day.

Our co-PI on the AFRL funded UV FEL project, Henry Helvajian, attended and gave an excellent presentation at this week's LPC Workshop, so we were able to use the opportunity to plan the next quarter of joint activities. We also were able to make some progress on our draft SOW for the AES CRADA on the 100 mA gun design with Alan Todd’s presence at the LPC Workshop.

WBS 4 (Injector):
The extractor cart drawings are ready for first review. The gun chamber was cleaned and is having its upstream and downstream tubes welded. Work continued on the new resistor effort.

Gun HVPS Tank - A meeting is scheduled this afternoon to review and approve the design of the upgrade HVPS Tank. All major components of the old Gun HVPS have been removed and stored in the Test Lab. The old HVPS Tank is ready for removal.

WBS 6 (RF):
Quarter Klystrons - The LCW manifold for the 100 kW klystron is in the Machine Shop for modifications. Two additional LCW manifolds will be constructed after the first one is approved. The RF Waveguide Loads were returned to the vendor for upgrading to 100 kW operations.

Quarter HVPS - In less than one day the old 3 phase variacs, transformers, and inductors were removed by a rigging company and the upgrade transformers (3000 lbs.) and inductors (1600 lbs.) were installed in the HVPS's. The schematic for the entire HVPS was reviewed, modified, and turned over to drafting. The installation of the other parts has begun.

Master Oscillator - The upgraded Master Oscillator is installed, connected to zones 3 & 4, and working fine. There are a few diagnostic features that will be added, but this task is basically complete.

WBS 8 (Instrumentation):
All of the control elements for the vacuum system of the IR upgrade were nailed down this week. Nomenclature was changed to maintain the naming standard as set up on the Main Machine and the IR Demo. Updated drawings are forthcoming.

Many manhours were spent moving and re-moving cables to support the necessary modifications to the cable trays in the vault. Three of the four changes required to remove interferences are complete.

New lamp rings for the beamviewer cameras are in the design phase and assembly of beamviewers continues.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o At Magnet Enterprises International in Oakland CA., they are winding a new first coil and have redone their potting mold to eliminate the problems. The potting of the new coil should complete next week during my visit. They are making a large effort in making cores and parts.
Injector Dipoles (DU/DV)
o Procurement is continuing via electronic methods. Many vendors have downloaded the web files and registered.
Arc Dipoles (GY, GX, GQ)
o Advanced Energy Systems (AES) analyzed the field integral information on the switch mode (to the UV Line) of the Bend Magnet (GX) along the half sagitta line. The field is not exactly half because of better permeability in the return legs. We will have to shunt current in these two magnets (in the half amp realm) to bring the field down, an easy solution to the problem. The field flatness and integral seem not to be grossly out of spec and detailed analysis is continuing. The vellums of all the drawings are ready for signature and the specifications are in checking.
o Copper was received by Wang NMR of CA for coil manufacturing. I will visit them next week.
Magnet Measurement
o Measurement of the dipole chamber’s magnetic properties started up using weld samples prepared by the company that welded the 180°chambers last time.
o Design is nearly complete for the probe system to measure absolute field of the Bend and Reverse Bend Dipoles.
3 inch quad (QX)
o We are waiting for the first batch OD quadrupoles to arrive.
Trim Quad (QT)
o All steel is rough machined and is in annealing. Winding forms are made. Mock up winding should start next week at Milhous Control of Virginia.
Sextupole (SF)
o In the arcs, we were anticipating providing vertical dipole correction field with some windings in the sextupoles. The 3D magnetic model of the sextupole at DULY Research shows that the superposed vertical dipole corrector field is coupled to the sextupole field. This is fatal to this concept, which needed a separate 10-pole corrector to correct the corrector. We will rather make a simple mu metal core corrector to do the task. There may be some beam line space problems that may force us to split the correction into two magnets, say on either side of the sextupole.
o DULY is now optimizing the field clamps on the sextupoleís magnetic model.
Octupole (OT)
o Work on this magnet is on furlough until the sextupole is designed.
Corrector Dipoles (DB, DJ, GC BH)
o No news on qualification of the new sets of Standard Corrector Dipoles (DB/DJ) because the measurement stand is set up to receive the quadrupoles after being fixed. Before that, the stand's problems prevented qualification of the magnets.
o We will now have to report on the new vertical corrector for the Arcs (GC) in this space as mentioned in Sextupole above.
o Chuck Hutton designed an IDEAs model of the single set of Haimson style correctors that is needed internal to the solenoid of the new gun configuration. The coils have to withstand 250 °C bake and have unique mounting requirements. Jeff Dale of the shop will make them using some very simple drawings from Chuck.
Beam Line and Vacuum
o AES is drawing up the chambers through out the arcs and is doing finite element analysis of the most problematic chamber, the chamber at the GX that has the optical cavity through it.
o Design work on girders and stands continued. Parts for the regions between the cryomodules were finally signed off.
o Work is delayed in the 6 quad telescope just before the first arc that was changed by David Douglas with his latest DIMAD revision. The original positions of the diagnostics (in this region, which was almost ready to sign off on,) were disturbed and we are waiting for Dave to consult about new positions. Dave’s presence was requested at Cornell workshop on ERL design this week.

WBS 11 (Optics):
The revised drawing package for the optical cavity assemblies has been released for checking. We received the enclosure we’ll use for testing the optical cavity stepper motors. We began working on the design for the insertable mirror assembly that is located immediately after the outcoupler assembly. We began tests with the Ga:In eutectic material we plan to use as a heat conductive medium for the edge-cooled outcoupler mirrors.

Meetings with possible vendors of drive laser components will be held at next week’s SPIE meeting.

Operations/Other Activities:
We worked with the team doing post analyses on the JTO-irradiated samples, clearing up some ambiguities with the nonmetallic samples. The team is now directing its attention on the metal samples. A presentation on the results was given at this week’s LPC User Workshop.